The speed of train A is 14mph slower than the speed of train B. Train A travels 190miles in the same time it takes train B to travel 260miles. What is the speed of each train?
Answer by edjones(8007)

Let b=the speed of b and a=speed of a.
b=a+14
t=time, s=speed, d=distance
t=d/s
d/s=d/s because the time of travel of each train is equal.
190/a=260/(a+14)
190(a+14)=260a
190a+2660=260a
70a=2660
a=38 mph
b=52 mph
.
Ed